An orchard has 864 orange trees. The number of rows exceeds the number of trees per row by 5 How many trees are there in each row

Answer by ikleyn(53751)

.
You have 2 (two) options:


Option 1

Solve the quadratic equation 

t*(t+5) = 864,

where t is the number of trees in each row.


Option 2

Find (MENTALLY) two factors of the number 864 that multiply up to 864 and differ by 5 units:

864 = 2*432 = 4*216 = 8*108 = 16*54 = 32*27 . . .    Oops !  You just found them:


Answer.  The number of trees in each row is 27.


Check.  27*(27+5) = 27*32 = 864.